Let f: Z®Z be such that f(x) = x +1.

Is f invertible? and if it is, what is its inverse?


1
Expert's answer
2021-03-29T08:58:42-0400

1) The map "f" is invertible if and only if "f" is bijective if and only if "f" is surjective and injective.

"f" is surjective because "\forall y \in \mathbb{Z} \exists x \in \mathbb{Z}: f(x) = x + 1 = y (x = y - 1)"

"f" is injective because "\forall x_1 \in \mathbb{Z}, \forall x_2 \in \mathbb{Z}, x_1 \neq x_2""f(x_1) = x_1 + 1 \neq x_2 + 1 = f(x_2)"

Thus, "f" is invertible.

2) The map "f ^{-1}: \mathbb{Z} \rightarrow \mathbb{Z}" is inverse to "f" if "f^{-1} f = f f^{-1} = e_{\mathbb{Z}}" , where "e_{\mathbb{Z}}" is the identical map "\mathbb{Z} \rightarrow \mathbb{Z}" .

"f^{-1}(y) = y - 1" because "f^{-1}f(x) = (x + 1) - 1 = x = \\ 
= (x - 1) + 1 = ff^{-1}(x)"